OpenStudy (anonymous):

Use pseudocode to specify a recursive algorithm to compute the nth value of the harmonic series, for some integer n.

OpenStudy (lukebluefive):

The harmonic series is defined by the following equation: \[\sum_{n=1}^{\infty} \frac{1}{n} = 1 + \frac{1}{2} + \frac{1}{3} + ...\] So in order to find the nth term, you simply keep adding elements until you reach the desired value of n. This is called your base case in recursion and it is critically important to get right. In this case, the base case would be like this: if (x == n) return x; Then, you would have to add an "else" statement that calls the harmonic function recursively, returning the sum of the next element with the elements added so far.